OpenStudy (anonymous):

In the course of one five-month period, the water level of a lake went down 6 in., up 12 in., down 3 in., down another 16 in., and up 4 in. How much had the lake level changed at the end of the five months?

OpenStudy (anonymous):

0-6+=-6+12=6-3=3-16=-13+4=...